OpenStudy (shamil98):

Which of the following statements are true? (more than one answer) F(x)=1/2 √x has the same domain and range as f(x) = cubed x. The graph of f(x) = cubed x / 2 looks like the graph of f(x) = cubed x, but will shrink it vertically by a factor of 1/2. The graph of f(x) = cubed x / 2 looks like the graph of f(x) = cubed x , but will shrink it horizontally by a factor of 1/2. The graph of f(x) = cubed x / 2 does not look like the graph of f(x) = cubed x because the roots are different.

OpenStudy (shamil98):

cubed x is just \[\huge \sqrt[3]{x}\]

ganeshie8 (ganeshie8):

not good wid horizontal shrink/stretch stuff but il try... take a snapshot and attach if psble :)

OpenStudy (shamil98):

Okay here it is

OpenStudy (shamil98):

I think its option B and D.

ganeshie8 (ganeshie8):

we can strike off first option as \(\sqrt{x} \) domain is only nonnegative. but \(\sqrt[3]{x}\) domain is all all reals

ganeshie8 (ganeshie8):

yes B is correct, 1/2 shrinks the graph vertically. let me see others :)

OpenStudy (shamil98):

okay.

ganeshie8 (ganeshie8):

yup B and D !

OpenStudy (shamil98):

Thank you :)